# Approaching Rate-Distortion Limits in Neural Compression with Lattice Transform Coding

### v1

@@ Abstract
Neural compression has brought tremendous progress in designing lossy compressors with good rate-distortion (RD) performance at low complexity. Thus far, neural compression design involves transforming the source to a latent vector, which is then rounded to integers and entropy coded. While this approach has been shown to be optimal in a one-shot sense on certain sources, we show that it is highly sub-optimal on i.i.d. sequences, and in fact always recovers scalar quantization of the original source sequence. We demonstrate that the sub-optimality is due to the choice of quantization scheme in the latent space, and not the transform design. By employing lattice quantization instead of scalar quantization in the latent space, we demonstrate that Lattice Transform Coding (LTC) is able to recover optimal vector quantization at various dimensions and approach the asymptotically-achievable rate-distortion function at reasonable complexity. On general vector sources, LTC improves upon standard neural compressors in one-shot coding performance. LTC also enables neural compressors that perform block coding on i.i.d. vector sources, which yields coding gain over optimal one-shot coding.
@@ 1 Introduction
where $\bm{y}=g_{a}(\bm{x})$, $\hat{\bm{y}}=Q(\bm{y})$, and $\hat{\bm{x}}=g_{s}(\hat{\bm{y}})$. NTC is considered a one-shot coding scheme of $\bm{x}$, because it compresses one source realization at a time.
@@
Figure 1: Nonlinear Transform Coding (NTC) recovers scalar quantization (ECSQ) on Gaussian i.i.d. sequences of any dimension $n$. Lattice Transform Coding (LTC) recovers optimal vector quantization (VQ) in each dimension, and approachces the rate-distortion function $R(D)$ without needing an exponential codebook search.
A recent line of work has investigated how close neural compressors perform in comparison to information-theoretic limits, which define the optimal performance any compressor may achieve. Yang and Mandt (2022) and Lei et al. (2022) estimate the rate-distortion function $R(D)$, which is the fundamental limit of lossy compression, for real-world data sources. Their results show that there exists a gap between NTC performance and $R(D)$. In general, the gap between NTC performance (or any one-shot code) and $R(D)$ can be attributed to (i) sub-optimality of NTC with regards to the optimal one-shot performance, and/or (ii) the lack of block coding (multi-shot coding), since $R(D)$ is achievable only under asymptotic block-lengths (Cover and Thomas, 2006).
The authors in (Wagner and Ballé, 2021; Bhadane et al., 2022) demonstrate that on stylized sources, NTC achieves the optimal one-shot coding performance. These sources are nominally high-dimensional, but have a low-dimensional representation, which we refer to as the latent source. The authors show that $g_{a},g_{s}$ can successfully extract the latent source, optimally quantize it, and map back to the original space, providing an optimal one-shot scheme as a whole.
@@
However, most of the sources they analyze have a one-dimensional uniform latent source $U\sim\mathrm{Unif}([a,b])$. While these are sufficient to analyze the role of $g_{a}$ and $g_{s}$, it does not provide insights on the role of the quantizer $Q$. This is because a uniform source $U$ can be optimally quantized using uniform scalar quantization, which is exactly what NTC uses in the latent space. For sources with higher-dimensional latent sources, where uniform scalar quantization is not necessarily optimal, it is unclear whether NTC can still provide optimal one-shot coding schemes.
An information-theoretic framework relevant to this question is that of companding (Bucklew, 1981). Companding defines a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, known as the compressor, which is applied to the source $\bm{x}\in\mathbb{R}^{n}$. $f(\bm{x})$ is then uniform lattice quantized, and passed to $f^{-1}$, known as the expander. Clearly, NTC can be thought of as an instantiation of a compander where $g_{a},g_{s}$ serve the role of $f,f^{-1}$. It is known that optimal companding is governed by the lattice’s packing efficiency and not $f$ (Linder et al., 1999). Uniform scalar quantization is equivalent to the integer lattice, which is sub-optimal in dimensions greater than 1.
This suggests that the results of one-shot optimality of NTC (Wagner and Ballé, 2021) may not hold in general when the latent source has dimensionality greater than 1. Surprisingly, this implies that when $\bm{x}$ contains an i.i.d. scalar sequence on its entries, NTC may not be able to compress it optimally. In this case, the latent source’s dimensionality is that of $\bm{x}$. The same implication holds for general vector sources, such as those from real-world data, where the latent source may reduce dimensionality but still have dimension greater than 1. In these cases, companding suggests the use of lattice quantization (LQ) in order to improve the one-shot coding performance of $\bm{x}$. Additionally, LQ provides a potential method to perform block coding of i.i.d. sequences of $\bm{x}$ in a scalable way, in order to bridge the gap between optimal one-shot coding and the rate-distortion function.
Figure 2: Nonlinear Transform Coding (Ballé et al., 2021). Lattice Transform Coding replaces scalar quantizer $Q$ with a lattice quantizer $Q_{\Lambda}$.
In this paper, we first provide evidence that NTC compresses such sources sub-optimally in Sec. 3. These results agree with those in the companding literature, and motivate the use of lattice quantization in the latent space as a potential way to improve performance, which we refer to as Lattice Transform Coding (LTC). LTC brings several challenges that require changes compared to NTC training, such as efficient lattice quantization algorithms, dealing with the non-differentiable quantization operation, integration of the entropy model over lattice regions, and accurate likelihood modelling. LTC is able to achieve optimal coding schemes for i.i.d. scalar sequences, while avoiding the exponential complexity of a direct codebook search under vector quantization (VQ); see Fig. 1. On general vector sources, LTC improves one-shot coding performance, as well as enable block coding of i.i.d. vector sequences. Our contributions are as follows.
@@
    We first demonstrate the inability of current NTC architectures to optimally compress i.i.d. scalar sequences. We show that this is due to the choice of scalar quantization in the latent space, and not the transform design.
@@
    We propose Lattice Transform Coding (LTC), and show that it is able to optimally compress i.i.d. sequences, matching the performance of optimal vector quantization, yet maintain reasonable complexity. We discuss various design choices regarding quantization, and entropy modelling that help achieve optimal performance. We demonstrate the coding gains achieved are due to the packing efficiency of the lattice in the latent space.
@@
    On general vector sources, we demonstrate LTC’s ability to improve one-shot coding schemes, as well as develop block coding schemes that compress i.i.d. blocks of vector sources, which are able to approach the rate-distortion bound.
@@ 2 Related Work
Rate-Distortion Theory.
Given a source $X\sim P_{X}$, the rate-distortion function (Cover and Thomas, 2006)
$R(D):=\min_{P_{\hat{X}|X}:\operatorname{\mathbb{E}}[\operatorname{\mathsf{d}}(X,\hat{X})]\leq D}I(X;\hat{X})$ (2)
describes the fundamental limit of compressing $X$. It is achievable in the limit of large block-lengths, where the rate and distortion are measured per-sample.
For a general vector source $\bm{x}$, its optimal one-shot coding scheme can be found via entropy-constrained vector quantization (ECVQ). ECVQ will approach $R(D)$ when $\bm{x}$ consists of i.i.d. sequences of $X$ and rate and distortion are divided by the sequence length. ECVQ directly enumerates quantization centers (codewords) in the space of $\bm{x}$, and minimizes an entropy-distortion objective,
@@
where $e(\bm{x}):=\operatorname*{arg\,min}_{\{\bm{c}_{i}\}_{i=1}^{\infty}}\operatorname{\mathsf{d}}(\bm{x},\bm{c}_{i})$. Since the solution is unknown in closed form, ECVQ is solved numerically by a generalized Lloyd’s algorithm (Chou et al., 1989).
@@
Although ECVQ is optimal, an issue that precludes its wider use is its high complexity; the number of codebook vectors required is exponential in the rate. This complexity issue is one reason why NTC is appealing, since scalar quantization in the transformed space is used, whose complexity is linear in the dimension of $\bm{y}$.
Rate-Distortion Estimation.
The work of Yang and Mandt (2022) and Lei et al. (2022) demonstrate a gap between NTC performance and estimates of $R(D)$ on real-world sources, and potentially attributes it to the lack of large block-length. We note that a comparison to $R(D)$ depends on the source and its relevant regimes on the rate-distortion trade-off. If an entire image is considered as a single sample, then typical regimes of interest include rates on the order of $10^{5}$ to $10^{6}$ bits-per-sample [^1] The bits-per-pixel (bpp) usually ranges between 0 and 2.. From Li and Gamal (2017), the gap between optimal one-shot coding and $R(D)$ is no more than $\log(R(D)+1)+5$. Thus, the gap becomes insignificant relative to the coding rate at such large rates; to compress a $256\times 256$ image patch at a bpp of 1, the gap corresponds to $\approx 0.00032$ bpp, which is much smaller than that reported in Yang and Mandt (2022). Therefore, for sources such as images, the performance gap can mostly be attributed to the sub-optimality of NTC as a one-shot coding scheme. For sources that have lower rate regimes of interest, the gap to $R(D)$ may be attributed to lack of large block-length as well.
Analysis of NTC.
The principles of nonlinear transform coding (NTC) are described in Ballé et al. (2021). An analysis of NTC was performed in Wagner and Ballé (2021); Bhadane et al. (2022), who consider synthetic sources and verify when NTC can recover the optimal (one-shot) rate-distortion performance. The focus is on determining when the analysis and synthesis transforms are able to recover the low-dimensional latent source; when this is possible, optimality is achieved empirically. However, the sources analyzed have a one-dimensional uniform latent source which uniform scalar quantization is optimal for; it is unknown whether sources with more complex latent sources can still be compressed optimally by NTC.
Vector Quantization in NTC.
@@
Although NTC is based on VAEs with scalar quantization in the latent space, several works have attempted to apply vector quantization in the latent space instead. Notable works include VQVAE (van den Oord et al., 2017), which performs a direct codebook search in the latent space and can be used as a compressor. PQVAE (El-Nouby et al., 2023) evaluates this approach with a product quantization scheme to improve scalability, demonstrating competitive performance on image compression benchmarks. Finally, NVTC (Feng et al., 2023) applies entropy-constrained vector quantization (ECVQ) algorithms (Chou et al., 1989) in the latent space, performing a direct codebook search. We show that LTC is able to recover the performance of ECVQ while avoiding a direct codebook search in the latent space, which is infeasible for large rates and dimensions.
Lattice Quantization.
@@
A lattice quantizer $Q_{\Lambda}$ (Conway and Sloane, 1999) consists of a countably infinite set of codebook vectors $\Lambda$ given by $\bm{u}\bm{G}$, where $\bm{G}\in\mathbb{R}^{n\times n}$ is the generator matrix of the lattice, and $\bm{u}$ are integer-valued row vectors. The Voronoi region of a lattice is given by $V(\bm{0}):=\{\bm{x}\in\mathbb{R}^{n}:Q_{\Lambda}(\bm{x})=\bm{0}\}$, where $Q_{\Lambda}(\bm{x}):=\operatorname*{arg\,min}_{\bm{v}\in\Lambda}\|\bm{x}-\bm{v}\|_{2}^{2}$ finds the closest lattice vector to $\bm{x}$ in $\Lambda$. In lattice theory, the objective is to find lattices that minimize the normalized second moment (NSM)

### v2

@@ Abstract
Neural compression has brought tremendous progress in designing lossy compressors with good rate-distortion (RD) performance at low complexity. Thus far, neural compression design involves transforming the source to a latent vector, which is then rounded to integers and entropy coded. While this approach has been shown to be optimal on a few specific sources, we show that it can be highly sub-optimal on synthetic sources whose intrinsic dimensionality is greater than one. With integer rounding in the latent space, the quantization regions induced by neural transformations, remain square-like and fail to match those of optimal vector quantization. We demonstrate that this phenomenon is due to the choice of scalar quantization in the latent space, and not the transform design. By employing lattice quantization instead, we propose Lattice Transform Coding (LTC) and show that it approximately recovers optimal vector quantization at reasonable complexity. On real-world sources, LTC improves upon standard neural compressors. LTC also provides a framework that can integrate structurally (near) optimal information-theoretic designs into lossy compression; examples include block coding, which yields coding gain over optimal one-shot coding and approaches the asymptotically-achievable rate-distortion function, as well as nested lattice quantization for low complexity fixed-rate coding.
@@ 1 Introduction
where $\bm{y}=g_{a}(\bm{x})$, $\hat{\bm{y}}=Q(\bm{y})$, $\hat{\bm{x}}=g_{s}(\hat{\bm{y}})$, and $\lambda>0$ controls the rate-distortion tradeoff. NTC is considered a one-shot coding scheme of $\bm{x}$, because it compresses one source realization at a time.
@@
Figure 1: Left: Nonlinear Transform Coding (NTC) recovers scalar quantization (ECSQ) on Gaussian i.i.d. sequences of any dimension $n$. Lattice Transform Coding (LTC) recovers optimal vector quantization (VQ) in each dimension, and approachces the rate-distortion function $R(D)$ without needing an exponential codebook search. Right: NTC architecture. LTC replaces scalar quantizer $Q$ with a lattice quantizer $Q_{\Lambda}$.
Optimal compression of $\bm{x}$ can be performed by vector quantization (VQ) Gersho & Gray (2012) albeit at high complexity which grows exponentially in rate and dimension. NTC, on the other hand, has low complexity. We are interested in understanding if NTC can recover VQ, and/or how its performance can be improved. If so, this would provide a low-complexity, source agnostic scheme capable of achieving the fundamental limits of compression.
@@
Wagner & Ballé (2021); Bhadane et al. (2022) demonstrate that on certain synthetic sources, NTC successfully achieves VQ performance. These sources are nominally high-dimensional, but have a low-dimensional representation, which we refer to as the latent source. The authors show that $g_{a},g_{s}$ can successfully extract the latent source, optimally quantize it, and map back to the original space, providing an optimal compression scheme as a whole. However, most of the sources they analyze have a one-dimensional uniform latent source $U\sim\mathrm{Unif}([a,b])$. While these are sufficient to analyze the role of $g_{a}$ and $g_{s}$, it does not provide insights on the role of the quantizer $Q$. This is because a uniform source $U$ can be optimally quantized using uniform scalar quantization (SQ), which is exactly what NTC uses in the latent space. For sources with higher-dimensional latent sources, where uniform SQ is not necessarily optimal, it is unclear whether NTC can still achieve optimality. Conceptually, given the high representation power of neural networks, one may hope that the transforms $g_{a}$ and $g_{s}$, in conjunction with the basic integer rounding in the latent space, can implement a more complex quantization such as VQ.
In this paper, we first provide evidence that on sources with higher latent dimensionality, NTC performs sub-optimally; the learned transforms are insufficient to overcome the sub-optimal latent SQ. Shown in Fig. 3, NTC quantization regions remain square-like (following integer rounding in the latent space), and do not match the hexagonal regions from VQ. Our observation is also aligned with the recent analytical approach of Shevchenko et al. (2023) for two-layer VAEs, showing fundamental incapability in recovering VQ for Gaussian sources. To improve performance, we propose to use lattice quantization (LQ), which provides improved space-packing efficiency over SQ, but reduced complexity compared to VQ. We refer to this framework (i.e., LQ in the latent space) as Lattice Transform Coding (LTC). On a variety of synthetic and real-world sources, we demonstrate how LTC improves upon NTC in terms of the rate-distortion trade-off, while avoiding the high complexity of a direct codebook search under VQ. For example, on isotropic Gaussian vectors, Fig. 1 shows how LTC achieves VQ performance without a codebook search.
The proposed framework of LTC further enables the integration of structurally-optimal solutions from information theory. In particular, we show how LTC enables design of (1) lower complexity solutions using nested lattices; and (2) near-optimal solutions using block coding from information theory. Overall, our experiments support LTC as a framework to achieve rate-distortion limits while mitigating high complexity.
Our contributions are as follows [^1] Code can be found at https://github.com/leieric/lattice-transform-coding..
@@
    We first demonstrate the inability of current NTC architectures to optimally compress sources with latent dimensionality larger than one; we conclude this is due to the choice of SQ in the latent space, and not the transform design.
@@
    We propose Lattice Transform Coding (LTC), a framework that replaces the latent integer rounding of NTC with general lattice quantization. LTC requires novel solutions for training, such as efficient lattice quantization algorithms, dealing with the non-differentiable quantization operation, integration of the entropy model over lattice regions, and accurate likelihood modeling.
@@
    LTC additionally enables integration of a variety of information-theoretic concepts into neural compression. Namely, we demonstrate how LTC allows the use of nested lattices in the latent space which provides a scheme with low encoding complexity, as well as implementation of block coding (necessary to achieve $R(D)$) which has applications in dataset compression.
  4. 4)
    We show the LTC framework can near-optimally compress interesting synthetic sequences for which we have a benchmark of optimality (e.g. i.i.d. sources, the Banana source), while maintaining reasonable complexity. Our results reveal the interplay between nonlinear transforms, entropy models and the choice of LQ.
  5. 5)
    For real-world sources with moderate to high dimensions, we demonstrate LTC’s ability to consistently improve upon the performance of NTC when the latent dimension is larger than one, and bridge up to 30% of the gap to $R(D)$.
@@ 2 Related Work
Rate-Distortion Theory. Given a source $X\sim P_{X}$, the rate-distortion function $R(D)$ describes the fundamental limit of compressing $X$ (Cover & Thomas, 2006). Compressing the source one sample at a time is known as one-shot coding. $R(D)$, on the other hand, is achievable asymptotically in the limit of large block-lengths when the source is compressed block by block and the rate and distortion are measured per-sample. For a general vector source $\bm{x}$, its optimal one-shot coding scheme can be found via entropy-constrained vector quantization (ECVQ).
ECVQ directly enumerates quantization centers (codewords) in the space of $\bm{x}$, and minimizes an entropy-distortion objective,
@@
where $e(\bm{x}):=\operatorname*{arg\,min}_{\{\bm{c}_{i}\}_{i=1}^{\infty}}\operatorname{\mathsf{d}}(\bm{x},\bm{c}_{i})$. Since the solution is unknown in closed form, ECVQ is solved numerically by a generalized Lloyd’s algorithm (Chou et al., 1989). Although ECVQ is optimal, an issue that precludes its wider use is its high complexity; the number of codebook vectors required is exponential in the rate and dimension. This complexity issue is one reason why NTC is appealing, since scalar quantization in the transformed space is used, whose complexity is linear in the dimension of $\bm{y}$. However, it is unknown whether NTC achieves optimality; this is an important question, since if NTC was optimal it would immediately offer a low-complexity solution that achieves ECVQ performance. Our work reveals that NTC is not optimal; however, LQ offers an optimal low-complexity solution for NTC.
Nonlinear Transform Coding. The principles of nonlinear transform coding (NTC) are described in Ballé et al. (2021).
@@
Although NTC is based on VAEs with SQ in the latent space, several works have attempted to apply VQ in the latent space instead. Notable works include VQVAE (van den Oord et al., 2017), which performs a direct codebook search in the latent space. PQVAE (El-Nouby et al., 2023) uses a product quantization scheme to improve scalability. Additionally, NVTC (Feng et al., 2023) applies ECVQ in the latent space, which also performs a direct codebook search. We show that LTC is able to recover the performance of ECVQ while avoiding a direct codebook search in the latent space, which is infeasible for large rates and dimensions.
@@
Lattice Quantization. A lattice quantizer $Q_{\Lambda}$ (Conway & Sloane, 1999) consists of a countably infinite set of codebook vectors $\Lambda$ given by $\bm{u}\bm{G}$, where $\bm{G}\in\mathbb{R}^{n\times n}$ is the generator matrix, and $\bm{u}$ are integer vectors. The Voronoi region of a lattice is $V(\bm{0}):=\{\bm{x}\in\mathbb{R}^{n}:Q_{\Lambda}(\bm{x})=\bm{0}\}$, where $Q_{\Lambda}(\bm{x}):=\operatorname*{arg\,min}_{\bm{v}\in\Lambda}\|\bm{x}-\bm{v}\|_{2}^{2}$ finds the closest lattice vector to $\bm{x}$ in $\Lambda$. In lattice theory, the objective is to find lattices that minimize $L_{\Lambda}$, the normalized second moment (NSM),
which measures the lattice’s packing efficiency. Recent work describing lattices which have the lowest known NSM can be found in (Agrell & Allen, 2023; Lyu et al., 2022; Agrell et al., 2024). Another objective is to find efficient quantization (decoding) algorithms (i.e., computing $Q_{\Lambda}(\bm{x})$), which requires solving the closest-vector problem (CVP); this is NP-hard in general. For many of the best-known lattices in dimensions up to 24 ($n=2$: hexagonal; $n=4$: $D_{n}^{*}$,; $n=8$: Gosset $E_{8}$; $n=16$: $\Lambda_{16}$ Barnes-Wall; and $n=24$: $\Lambda_{24}$ Leech) efficient CVP solvers exist (Conway & Sloane, 1999).
Classically, lattices are used to implement both fixed and variable-rate quantizers (Zamir et al., 2014). For the former, nested lattices shape the used part of the lattice and create a finite codebook with index-based encoding. For the latter, the entire (infinite) lattice is used, but entropy coding is applied; this is known as entropy-constrained lattice quantization (ECLQ). Fixed-rate quantization trades lower encoding complexity for rate-distortion performance. For both cases, in the large rate regime, LQ achieves optimality when the dimension is asymptotically large. For example, for ECLQ,
$\lim_{D\rightarrow 0}\frac{1}{n}\left(H_{\text{ECLQ}}(D)-R_{\bm{x}}(D)\right)$ $=\frac{1}{2}\log(2\pi eL_{\Lambda}),$ (3)
where $R_{\bm{x}}(D)$ is the rate-distortion function of $\bm{x}$ and $H_{\text{ECLQ}}(D)$ is the rate of ECLQ at distortion $D$ (Zamir et al., 2014, Ch. 5). If $G_{n}$ is the minimal NSM of any lattice in $n$ dimensions, $\lim_{n\rightarrow\infty}G_{n}={1}/({2\pi e})$, and thus (3) converges to 0. For the integer lattice (SQ), the NSM is $1/12$, which yields a gap of $\approx 0.255$ bits per dimension. Thus, LQ offers the maximal benefit when the rate-per-dimension is relatively small. While this may seem limiting, many real-world sources operate at low rate-per-dimension. For example, in NTC-based image compression, the rate-per-latent dimension is around the same as the bits-per-pixel used, which is typically between 0 and 2. While classical transform coding may not have benefited much from LQ since the transform domain of images (e.g, DCT) is typically sparse, this is not true in NTC where the rate is spread more evenly across all dimensions (Cheng et al., 2020; Bhadane et al., 2021) resulting in a lower bit-per-dimension for the quantizer.
Companding. Companding defines a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, known as the compressor, which is applied to the source $\bm{x}\in\mathbb{R}^{n}$. $f(\bm{x})$ is then uniform lattice quantized, and passed to $f^{-1}$. Clearly, NTC can be thought of as an instantiation of a compander where $g_{a},g_{s}$ serve the role of $f,f^{-1}$. Companding has typically been analyzed in the (asymptotically) large rate, small distortion regime. Optimal scalar quantization can be implemented using scalar companding (Bennett, 1948). For the vector case, vector companders cannot implement optimal VQ (Gersho, 1979; Bucklew, 1981; 1983), even when the best lattices are used, except for a very small class of sources. Despite this, it is still possible optimal companding may perform near-optimal VQ. This is supported by (Linder et al., 1999) which has a similar asymptotic result as (3). Due to this, we seek to utilize them in our learning-based framework and learn the companding function as guided by the companding literature.
Lattice Quantization and Neural Compression. The first work to integrate a lattice quantizer with NTC was LVQAC (Zhang & Wu, 2023), which proposed the use of the $D_{n}^{*}$ lattice in the latent space, and showed some performance improvement over scalar quantization. However, LVQAC does not generalize to other lattices, and the $D_{n}^{*}$ is known to perform poorly compared to many other known lattices. In contrast, our approach is generalizable to high-performing lattices, and we further show how the performance of learned compressors depend on the space-packing efficiency of the lattice quantizer. LVQ-VAE (Kudo et al., 2023), a contemporary work to ours, seeks to exploit correlations among latent features by extending the hyperprior entropy model (Ballé et al., 2018) with a covariance matrix along quantized vectors. To mitigate the costly search over vector codewords, Kudo et al. (2023) uses lattice quantization in the latent space in a manner similar to our work. However, their work does not shed light on the role of lattice quantization in NTC. Moreover, their approach suffers from high complexity. In contrast, we seek to understand the precise role that quantization plays in NTC, and lattice quantization emerges as a solution to achieve optimal rate-distortion performance. We additionally show how nested lattices can help achieve optimality at much lower complexity. Finally, Cao et al. (2024) is another contemporary paper that provides a low-complexity lattice VQ solution for image compression; however, details are sparse sin their manuscript for comparison.
3 NTC is Sub-Optimal
Let $\bm{x}\sim P_{\bm{x}}$ be the vector source. We train an NTC model following Ballé et al. (2021) using a factorized entropy model for the distribution of the latent $\bm{y}$. We use the objective in (1), with $\operatorname{\mathsf{d}}(\bm{x},\hat{\bm{x}})=\|\bm{x}-\hat{\bm{x}}\|_{2}^{2}$, and compare with (optimal) entropy-constrained vector quantization (ECVQ). We first consider the case when the $n$ entries of $\bm{x}$ are i.i.d. with marginal $S\sim\mathcal{N}(0,1)$. Shown in Fig. 1, NTC with $n=1$ is able to recover ECVQ with $n=1$, also known as entropy-constrained scalar quantization (ECSQ). This is perhaps expected, since the task of the transforms is to scale the 1-D Gaussian to be optimally quantized on an integer grid in 1-D, and then scale the quantized latent to form the reconstruction. Since the source is intrinsically one-dimensional, scalar quantization appears to be sufficient in the latent space to implement optimal quantization.
For $n=2$, however, NTC is unable to recover ECVQ, and always achieves the same performance as that of ECSQ.
Figure 2: Rate-distortion, 2-D Gaussians.
Figure 3: Quantizer regions. Left: NTC. Right: ECVQ. Top: 2-D Gaussian source. Bottom: Banana source.
Shown in Fig. 3, we see that the rate-distortion performance for ECVQ at $n=2$ is superior to that of ECSQ. However, NTC at $n=2$ merely recovers ECSQ performance, and fails to outperform it. Visualizing the quantization regions on the 2-D Gaussian source (Fig. 3) reveals why. NTC is merely implementing a scalar quantizer of the source, perhaps with a rotation, which is equivalent to applying ECSQ to the entries of $\bm{x}$ independently. In 2-D, this induces square quantization regions in the source space. In contrast, the optimal ECVQ in $n=2$ has hexagonal regions, not square. A similar phenomenon occurs for the Banana source (Ballé et al., 2021). It is well-known that hexagonal tessellations yield superior packing efficiency compared to square tessellations (Conway & Sloane, 1999). Thus, in 2-D, for NTC to be optimal, the quantization regions (induced by latent SQ and learned transforms) in the source space need to be hexagonal.
We found that when the latent space is rounded to integers, varying the depth, width, nonlinearities, biases, and even increasing the latent dimension to a higher-dimensional space did not help; all resulted in square regions in the source space. For the i.i.d. Gaussian case, setting $n>2$ did not help either; all models recovered the performance of ECSQ at best. This surprising observation reveals that the learned transforms $g_{a}$ and $g_{s}$ seem unable to overcome use of scalar quantization. Moreover, it confirms that when the latent dimensionality of the source is greater than one, using SQ in the latent space, which is optimal for uniform scalar sources, does not result in optimality.
Several related works in the literature appear to support this observation. Shevchenko et al. (2023) analyzes 2-layer VAEs with a form of scalar quantization in the latent space, and show that the optimal performance is given by SQ, not VQ. Results in the companding literature (Gersho, 1979; Linder et al., 1999) additionally suggest that a compressor such as NTC is (i) unable to perform exact VQ for most sources, and that (ii) the performance is determined by the choice of the lattice in the latent space, which for NTC, is equivalent to the (sub-optimal) integer lattice.
To achieve VQ performance, these observations suggest that learned transforms are insufficient in overcoming sub-optimal quantization in the latent space, and that improving said latent quantization is required to achieve optimality. As will be shown, the proposed LTC is able to recover ECVQ regions and performance, with increasing gains over NTC as dimensionality increases (Fig. 1).
4 Lattice Transform Coding
We seek a compressor that can provide better performance than scalar quantization yet avoid the high complexity of vector quantization. Using a lattice quantizer $Q_{\Lambda}$ provides an intermediate solution that manages both complexity and space-filling efficiency. It generalizes scalar quantization, which is equivalent to the integer lattice. It is known that many lattices can provide far more space-filling efficiency compared to the integer lattice. We first discuss how to extend current NTC designs to integrate the use of lattices. We then discuss a method to reduce the entropy coding complexity of LTC by using nested lattice quantization. Finally, we discuss how LTC enables one to approach the asymptotic rate-distortion function $R(D)$ for any source.
4.1 From Scalar to Lattice Quantization
The primary challenges when integrating lattice quantization with (1) include computing the quantization itself and computing integrals over the lattice cell for likelihood modelling.
Performing lattice quantization. The forward-pass of the lattice quantizer $Q_{\Lambda}$ requires the CVP algorithm of the lattice; exact algorithms we use for the lattices in this paper are described in Sec. A.1.
To allow differentiation for the backwards pass, we follow standard practices in NTC and use either the straight-through estimator (STE) or dithered quantization with uniform noise over the lattice cell. The STE approach can use (1) directly as an objective during training, but may lead to training instabilities. The uniform noise proxy has stabler training, but requires a rate-distortion proxy loss of (1) using a connection to dithered quantization (Ballé et al., 2021). We also assume that the lattice’s generator matrix $\bm{G}$ has been normalized to be unit volume, i.e. $\mathrm{Vol}(V(\bm{0}))=\sqrt{\det(\bm{G}\bm{G}^{\top})}=1$.
We use STE for the distortion term $\operatorname{\mathbb{E}}_{\bm{x}}[\operatorname{\mathsf{d}}(\bm{x},g_{s}(Q_{\Lambda}(g_{a}(\bm{x}))))]$. Specifically, we compute the quantized output $\hat{\bm{y}}=Q_{\Lambda}(g_{a}(\bm{x}))$ during the forward pass via
The rate term, however, requires more care.